Shutdown tylansige 2001-06-04 16:34:33: Heaters not working

seymour at snf.stanford.edu seymour at snf.stanford.edu
Tue Jun 5 10:38:10 PDT 2001


removed and cleaned contacts on the temp control unit board. Checked over temp operation. restarted pumpsige. temp now at  +500 degs all operations look ok. 




More information about the tylansige-pcs mailing list